OpenStudy (anonymous):

Simplify the expression (x^3y^-6/x^6y^6)^1/3 @StudyGurl14

OpenStudy (anonymous):

1x^-1

OpenStudy (anonymous):

How did you get that answer?

OpenStudy (anonymous):

It's hard to explain but i'll try to explain it to you ;)

OpenStudy (anonymous):

Thanks for your effort. :)

OpenStudy (anonymous):

\[The problem \to solve is: (x^3y^-6/x^6y^6)^1/3 Multiply the exponent of y by 6giving The answer is y^6 evaluates \to Multiply the exponent of x by 6giving The answer is x^6 evaluates \to The 1 becomes a 1/1 The exponent, -6, is negative, so we'\ll have \to handle \it a little bit differently here. But for now, we'\ll treat \it as \it was positive, and handle \it \in the usual manner... that is, as if \it were at +6. Multiply the exponent of y by 6giving Now, \because of the negative exponent, we take this result, and turn \it upside-down. That is, a number becomes 1/the number, or a fraction is literally flipped upside-down. This effectively "flips" the entire result upside down, but that's exactly what negative exponents do! In this case... The 1 becomes a 1/1 y^6 becomes 1/y^6. The answer is y^-6 evaluates \to Multiply the exponent of x by 3giving The answer is x^3 evaluates \to Multiply and Multiply the and Multiply and The just \gets copied along. The just \gets copied along. The answer is x^3*y^-6 evaluates \to To divide by Divide by by subtracting the exponents, and keeping the x, \to get The just \gets copied along \in the numerator. The answer is x^3*y^-6/x^6 evaluates \to Multiply and Multiply the and Multiply and Combine the and by adding the exponents, and keeping the y, \to get 1 The just \gets copied along. The answer is x^3*y^-6/x^6*y^6 evaluates \to (x^3*y^-6/x^6*y^6)^1/3 evaluates \to \]

OpenStudy (anonymous):

haha i tried to insert it all it aint work i'll take a pic for you then

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

ok i got the pics :)

OpenStudy (anonymous):

there you go :)

rishavraj (rishavraj):

if its \[(\frac{ x^3 y ^{-6} }{ x^6 y^6 })^{\frac{ 1 }{ 3}}\] then asnwer is \[\frac{ 1 }{ x y^4 } = x ^{-1} y ^{-4}\]

OpenStudy (anonymous):

may i ask how do you get the y-4?

OpenStudy (anonymous):

y^-4*

rishavraj (rishavraj):

@Najia2000 thts the answer .....:)